40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 7 19:42:01 2018, gautam, Update, General, IFO checkup and DRMI locking prep 
    Reply  Wed Nov 7 23:07:11 2018, gautam, Update, LSC, DRMI locking recovered Screenshot_from_2018-11-07_23-05-58.pngSRCL2MICH_crosscpl.pdfPRCangularCoh_rot.pdf
       Reply  Sat Nov 10 17:32:33 2018, gautam, Update, LSC, Nulling MICH->PRCL and MICH->SRCL sensMat_2018-11-10.pdf
          Reply  Tue Nov 13 18:09:24 2018, gautam, Update, LSC, Investigation of SRCL-->MICH coupling TF_sweptSineMeas.pdfdigitalDemod.pdfdigitalDemod_hist.pdfDRMI_LSCspectrogram.pdf
             Reply  Fri Nov 16 00:47:43 2018, gautam, Update, LSC, More DRMI characterization PRCff.pdfDRMI_withPRCff.png
Message ID: 14285     Entry time: Wed Nov 7 23:07:11 2018     In reply to: 14284     Reply to this: 14288
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI locking recovered 

I had some success today. I hope that the tweaks I made will allow working with the DRMI during the day as well, though it looks like the main limiting factor in lock duty cycle is angular stability of the PRC.

  • Since there has been some change in the light levels / in vacuum optical paths, I decided to be a bit more systematic.
  • Initial guess of locking gains / demod phases was what I had last year.
  • Then I misaligned SRM, and locked PRMI, for the sideband resonant in the PRC (but still no arm cavities, and using 1f Refl error signals).
  • Measured loop TFs, adjusted gains, re-enabled boosts.
  • Brought the SRM back into the picture. Decided to trigger SRCL loop on AS110I rather than the existing POP22I (because why should 2f1 signal buildup carry information about SRCL?). New settings saved to the configure script. Reduced MICH gain to account for the SRC cavity gain.
  • Re-measured loop TFs, re-adjusted gains. More analysis about the state of the loops tomorrow, but all loops have UGF ~100-120 Hz.
  • Ran some sensing lines - need to check my sensing matrix making script, and once I get the matrix elements, I can correct the error signal demod phasing as necessary.

[Attachment #1]: Repeatable and reliable DRMI locks tonight, stability is mainly limited by angular glitches - I'm not sure yet if these are due to a suspect Oplev servo on the PRM, or if they're because of the tip-tilt PR2/PR3/SR2/SR3.

[Attachment #2]: A pass at measuring the TF from SRCL error point to MICH error point via control noise re-injection. I was trying to measure down to 40 Hz, but lost the lock, and am calling it for the night.

[Attachment #3]: Coherence between PRM oplev error point and beam spot motion on POP QPD.

Note that the MICH actuation is not necessarily optimally de-coupled by actuating on the PRM and SRM yet (i.e. the latter two elements of the LSC output matrix are not precisely tuned yet).

What is the correct way to make feedforward filters for this application? Swept-sine transfer function measurement? Or drive broadband noise at the SRCL error point and then do time-domain Wiener filter construction using SRCL error as the witness and MICH error as the target? Or some other technique? Does this even count as "feedforward" since the sensor is not truly "outside" the loop?

Attachment 1: Screenshot_from_2018-11-07_23-05-58.png  43 kB  | Hide | Hide all
Screenshot_from_2018-11-07_23-05-58.png
Attachment 2: SRCL2MICH_crosscpl.pdf  5 kB  | Hide | Hide all
SRCL2MICH_crosscpl.pdf
Attachment 3: PRCangularCoh_rot.pdf  62 kB  | Hide | Hide all
PRCangularCoh_rot.pdf
ELOG V3.1.3-